OpenStudy (anonymous):

use the exponent laws to simplify the following until you have only a single power. then evaluate.

OpenStudy (anonymous):

\[2^{4} \times(4^{5}\div4^{2}) \over (6 + 7^{-3})^{0}\]

OpenStudy (anonymous):

ok sarah, first, (4^5)/(4^2)=4^3 4^3=2^4, so we have (2^4)*(2^4) in the denominator, but i can't see the numenator clearly, can u type it again

OpenStudy (anonymous):

i dont even know what demoninator and numerator mean

OpenStudy (anonymous):